# Select and Distill: Selective Dual-Teacher Knowledge Transfer for Continual Learning on Vision-Language Models

### v1

@@ Abstract
Large-scale vision-language models (VLMs) have shown a strong zero-shot generalization capability on unseen-domain data. However, when adapting pre-trained VLMs to a sequence of downstream tasks, they are prone to forgetting previously learned knowledge and degrade their zero-shot classification capability. To tackle this problem, we propose a unique Selective Dual-Teacher Knowledge Transfer framework that leverages the most recent fine-tuned and the original pre-trained VLMs as dual teachers to preserve the previously learned knowledge and zero-shot capabilities, respectively. With only access to an unlabeled reference dataset, our proposed framework performs a selective knowledge distillation mechanism by measuring the feature discrepancy from the dual teacher VLMs. Consequently, our selective dual-teacher knowledge distillation would mitigate catastrophic forgetting of previously learned knowledge while preserving the zero-shot capabilities from pre-trained VLMs. Through extensive experiments on benchmark datasets, we show that our proposed framework is favorable against state-of-the-art continual learning approaches for preventing catastrophic forgetting and zero-shot degradation.
@@ 1 Introduction
To address the degradation of zero-shot capabilities during sequentially model fine-tuning, very recent work ZSCL [47] is proposed to regularize the optimization on the current task through the guidance of the original pre-trained VLMs (e.g., CLIP [33]). More specifically, ZSCL [47] distills the knowledge from a teacher VLM, which is frozen without fine-tuning, to constrain the fine-tuned student VLM using an unlabeled reference dataset (e.g., ImageNet [11]). In this way, the student VLM could preserve the intrinsic capability of zero-shot transfer in VLMs during fine-tuning without requiring the assessment of the pre-training dataset. However, such a manner solely considers the zero-shot capability of VLMs. The knowledge learned from each previous task cannot readily be preserved when only unlabeled reference dataset accessible, since the sampled reference images might have dissimilar appearance or semantics to past tasks. Therefore, incrementally expanding the learned capability from previous tasks remains challenging yet unsolved.
@@ 3.2 Selective Dual-Teacher Knowledge Transfer on VLMs
Algorithm 1 Selective Dual-Teacher Knowledge Transfer
Output: The final fine-tuned model $g_{K}$.
1:  for $k$ in $1:K$ do
2:     Freeze $g_{0}$ as the pre-trained knowledge teacher.
3:     Freeze $g_{k-1}$ as the previously fine-tuned knowledge teacher.
4:     Initialize the current model $g_{k}$ by $g_{k-1}$.
5:     for $e$ in $E$ do
6:        while not traverse over all current data $\mathcal{T}^{k}$ do
7:           Sample a batch of current data $B^{k}$.
8:           Sample a batch of ref data $B^{\text{ref}}$.
9:           Calculate $\mathcal{L}_{\text{CE}}$ with the current data $B^{k}$.
10:           Calculate Eq. 3 with $g_{0}$, $g_{k-1}$, and $B^{\text{ref}}$.
11:           Update $g_{k}$ with loss function Eq. 4.
12:        end while
13:     end for
14:  end for
@@ 3.3 Training and Inference
We summarize the training algorithm of our proposed Selective Dual-Teacher Knowledge Transfer in Algorithm 1. Note that at each stage $k$, we do not have access to data from previous tasks $\{\mathcal{T}^{1},\cdots,\mathcal{T}^{k-1}\}$. After sequentially fine-tuning over all $K$ different tasks, we derive a final model $g_{K}$ that exhibits zero-shot classification capabilities with catastrophic forgetting suppressed.
@@ 4.3 Evaluation Protocol
Specifically, we pick up the first training sequence $\mathcal{S}^{1}$ in the following order: (Aircraft, DTD, EuroSAT, Flowers, Food, Pets, Cars, UCF101), and the next $7$ training sequences can be derived through Eq. 6.
@@ 4.4 Main Result
We compare our proposed framework with several baseline methods. Continual FT is the most straightforward strategy that continually fine-tunes the model to the current task. LwF [23] proposes to distill knowledge from the previous model with the current data. iCaRL [34] maintains a memory buffer that stores previous data and performs knowledge distillation to acquire knowledge from the previous models. ZSCL [47] is the most related method to our approach, which also introduces a reference dataset and distills knowledge from the pre-trained CLIP model. In addition, they further apply a weight-space ensemble for certain intervals to ensure a gradual transition of model parameters.
@@
Method /// Sequence $\mathcal{S}^{1}$ $\mathcal{S}^{2}$ $\mathcal{S}^{3}$ $\mathcal{S}^{4}$ $\mathcal{S}^{5}$ $\mathcal{S}^{6}$ $\mathcal{S}^{7}$ $\mathcal{S}^{8}$ Catastrophic forgetting (normal-$\downarrow$) Continual FT 10.98 10.60 8.80 19.17 10.11 11.95 15.19 9.48 LwF [23] 10.38 6.52 6.37 10.22 7.99 7.70 10.41 8.91 iCaRL [34] 8.42 7.00 6.45 10.21 7.03 7.33 9.68 8.23 ZSCL [47] 4.67 2.35 2.13 2.97 3.15 4.28 4.89 4.70 Ours 1.70 1.16 0.89 1.04 0.59 1.34 1.12 1.79 Zero-shot degradation (normal-$\downarrow$) Continual FT 24.81 23.58 19.54 16.46 22.22 19.02 19.54 24.02 LwF [23] 10.75 10.23 8.63 8.25 12.02 10.33 8.98 11.01 iCaRL [34] 13.77 12.68 11.28 12.14 13.20 13.20 13.09 14.01 ZSCL [47] 3.44 3.94 4.02 2.85 3.79 2.31 1.86 1.84 Ours 1.55 2.04 1.21 1.92 2.79 2.18 1.90 2.08 Average accuracy (normal-$\uparrow$) Continual FT 76.16 76.24 78.03 68.69 76.64 75.44 72.71 77.45 LwF [23] 76.78 80.45 80.65 77.52 79.64 79.45 77.31 78.70 iCaRL [34] 77.99 79.77 79.93 76.66 79.26 79.08 77.06 78.61 ZSCL [47] 81.89 83.98 84.30 83.49 83.41 82.38 81.92 81.97 Ours 84.48 84.92 84.97 84.89 85.50 85.07 85.02 84.52
@@
Method /// Sequence $\mathcal{S}^{1}$ $\mathcal{S}^{2}$ $\mathcal{S}^{3}$ $\mathcal{S}^{4}$ $\mathcal{S}^{5}$ $\mathcal{S}^{6}$ $\mathcal{S}^{7}$ $\mathcal{S}^{8}$ Catastrophic forgetting (normal-$\downarrow$) Continual FT 11.17 10.89 10.16 20.12 10.57 12.14 15.62 9.80 LwF [23] 9.56 6.38 6.93 11.09 8.37 7.69 10.24 8.44 iCaRL [34] 8.43 6.90 6.83 10.69 7.09 7.37 10.17 8.66 ZSCL [47] 4.21 1.41 2.08 3.32 2.85 4.39 5.22 5.13 Ours 1.92 1.53 0.97 1.14 0.58 1.55 1.29 1.81 Zero-shot degradation (normal-$\downarrow$) Continual FT 24.54 24.10 19.53 17.60 21.96 18.92 20.26 24.31 LwF [23] 11.94 11.82 8.27 9.99 13.36 11.47 10.95 12.63 iCaRL [34] 13.02 12.78 10.89 11.81 12.74 12.87 11.92 13.34 ZSCL [47] 3.59 4.71 4.17 2.81 3.55 1.97 1.47 2.30 Ours 1.44 1.80 1.01 1.53 2.17 1.80 1.65 1.82 Average accuracy (normal-$\uparrow$) Continual FT 75.17 75.13 76.01 67.17 75.54 74.47 71.66 76.40 LwF [23] 74.14 77.44 77.94 74.89 77.30 77.43 75.50 76.51 iCaRL 76.97 78.82 78.57 75.43 78.08 78.10 75.70 77.52 ZSCL [47] 80.49 82.54 82.99 82.08 82.17 80.99 80.30 80.09 Ours 83.35 83.57 83.88 83.70 84.46 83.82 83.89 83.43
@@
As shown in Tab. 1, we present the quantitative comparisons with different methods on the MTIL benchmark [47]. The results demonstrate that our method outperforms SOTA CL approaches, showing the effectiveness of our proposed method. By leveraging dual teachers with the proposed teacher selective mechanism, our framework is capable of alleviating catastrophic forgetting on all training sequences with less than $2\%$ of performance degradation. In addition, the zero-shot classification capability can be properly preserved by our proposed learning framework.
@@ 4.5 Analysis
The performance on the first dataset in a training sequence inevitably suffers from the most severe catastrophic forgetting during continual learning. In Fig. 4, we plot the evaluation results on the first task through all training rounds to visualize the degree of catastrophic forgetting. The results clearly demonstrate that our method effectively maintains stable performance on the initial task even after multiple training rounds, verifying the effectiveness of our method in alleviating catastrophic forgetting on previous tasks. More experimental results are provided in our supplementary material.
@@
In addition to maintaining the knowledge learned from previous tasks, it is also crucial to preserve the zero-shot transferability for continual learning on VLMs. Intuitively, the performance of the last dataset in the training sequence experiences the most significant zero-shot degradation before fine-tuning on this dataset. Fig. 5 indicates the level of zero-shot degradation for the last dataset. The results show that we are able to keep almost the same zero-shot performance compared with the original pre-trained CLIP, demonstrating the effectiveness of our method in preserving the pre-trained zero-shot classification capability during sequential fine-tuning. We performed additional experiments on different training sequences in our supplementary material.
@@
In Sec. 3.2, we argue that the dual-teacher discrepancy is highly related to determining whether an image is visually similar to previously fine-tuned data. Here, we verify the idea with empirical analysis. At stage $k=2$, given a model $g_{1}$ fine-tuned on the Aircraft dataset ($\mathcal{T}^{1}$) and a pre-trained model $g_{0}$, we calculate the average dual-teacher discrepancy between $g_{1}$ and $g_{0}$ across various datasets, as shown in Tab. 3. The results indicate that the average dual-teacher discrepancy $d$ on the Aircraft dataset ($\mathcal{T}^{1}$) is significantly greater than those on other datasets that have not been fine-tuned by $g_{1}$. This observation suggests that previously fine-tuned data (i.e., Aircraft) does have a larger dual-teacher discrepancy, whereas data that does not align with $\mathcal{X}^{1}$ shows a notably lower discrepancy, empirically verifying our intuition in Eq. 1.
@@
  • After the model has been sequentially trained on multiple rounds (with the last task as Cars), our scheme is still able to select reference images closer to prior tasks (e.g., Aircraft), as visualized in the right part of Fig. 6.
@@ 5 Conclusion
In this paper, we proposed a Selective Dual-Teacher Knowledge Transfer framework for continual learning, which uniquely tackles catastrophic forgetting and preserve the model’s zero-shot generalization ability simultaneously. By leveraging the most recent fine-tuned and the original pre-trained VLMs as dual teachers, our framework performs selective knowledge distillation from the above models based on the proposed dual-teacher discrepancy. Such discrepancy is simply observed from a irrelevant reference dataset and does not require any label supervision during the distillation process. We conducted comprehensive experiments, including comparisons to state-of-the-art continual learning methods with extensive analysis. We quantitatively and qualitatively verified the effectiveness of our proposed framework over existing approaches.

### v2

@@ Abstract
Large-scale vision-language models (VLMs) have shown a strong zero-shot generalization capability on unseen-domain data. However, adapting pre-trained VLMs to a sequence of downstream tasks often leads to the forgetting of previously learned knowledge and a reduction in zero-shot classification performance. To tackle this problem, we propose a unique Selective Dual-Teacher Knowledge Transfer framework that leverages the most recent fine-tuned and the original pre-trained VLMs as dual teachers to preserve the previously learned knowledge and zero-shot capabilities, respectively. With only access to an unlabeled reference dataset, our proposed framework performs a selective knowledge distillation mechanism by measuring the feature discrepancy from the dual-teacher VLMs. Consequently, our selective dual-teacher knowledge distillation mitigates catastrophic forgetting of previously learned knowledge while preserving the zero-shot capabilities of pre-trained VLMs. Extensive experiments on benchmark datasets demonstrate that our framework is favorable against state-of-the-art continual learning approaches for preventing catastrophic forgetting and zero-shot degradation. Project page: https://chuyu.org/research/snd.
@@ 1 Introduction
To address the degradation of zero-shot capabilities during sequential model fine-tuning, very recent work ZSCL [50] is proposed to regularize the optimization on the current task through guidance from the original pre-trained VLMs (e.g., CLIP [34]). More specifically, ZSCL [50] distills the knowledge from a teacher VLM, which remains frozen without fine-tuning, to constrain the fine-tuned student VLM using an unlabeled reference dataset (e.g., ImageNet [11]). This approach allows the student VLM to preserve the intrinsic zero-shot transfer capability of VLMs during fine-tuning without requiring the assessment of the pre-trained dataset. However, such a manner solely considers the zero-shot capability of VLMs. The knowledge learned from previous tasks cannot be readily preserved since the pre-trained model has never fine-tuned on previous tasks, resulting in limited performance improvement against catastrophic forgetting. Therefore, incrementally expanding the learned capability from previous tasks remains a challenging and unsolved problem.
@@ 3.3 Training and Inference
Following previous settings for continual learning on Vision-Language Models [50], we fine-tune the original pre-trained model CLIP [34] to the downstream tasks in a sequential manner. We summarize the training algorithm in our supplementary material. Note that at each stage $k$, we do not have access to data from previous tasks $\{\mathcal{T}^{1},\cdots,\mathcal{T}^{k-1}\}$. After sequentially fine-tuning over all $K$ different tasks, we derive a final model $g_{K}$ that exhibits zero-shot classification capabilities with catastrophic forgetting suppressed.
@@ 4.3 Evaluation Protocol
Specifically, we pick up the first training sequence $\mathcal{S}^{1}$ in the following order: (Aircraft, DTD, EuroSAT, Flowers, Food, Pets, Cars, UCF101), and the next $7$ training sequences can be derived through Eq. 6. We leave the detailed order of tasks for each training sequence in the supplementary.
@@ 4.4 Main Result
We compare our proposed framework with several baseline methods. Continual FT is the most straightforward strategy that continually fine-tunes the model to the current task. LwF [24] proposes to distill knowledge from the previous model with the current data. iCaRL [35] maintains a memory buffer that stores previous data and performs knowledge distillation to acquire knowledge from the previous models. ZSCL [50] is the most related method to our approach, which also introduces a reference dataset and distills knowledge from the pre-trained CLIP model. In addition, they further apply a weight-space ensemble for certain intervals to ensure a gradual transition of model parameters. MoE-Adapters [48] is the newest state-of-the-art that involves incremental adapters as mixture-of-experts [18, 38] upon a frozen CLIP model, and further apply an automatic selector to allocate data to the experts during test phase.
@@
Method /// Sequence $\mathcal{S}^{1}$ $\mathcal{S}^{2}$ $\mathcal{S}^{3}$ $\mathcal{S}^{4}$ $\mathcal{S}^{5}$ $\mathcal{S}^{6}$ $\mathcal{S}^{7}$ $\mathcal{S}^{8}$ Mean Catastrophic forgetting ($\downarrow$) Continual FT 10.98 10.60 8.80 19.17 10.11 11.95 15.19 9.48 12.04 LwF [24] 10.38 6.52 6.37 10.22 7.99 7.70 10.41 8.91 8.56 iCaRL [35] 8.42 7.00 6.45 10.21 7.03 7.33 9.68 8.23 8.04 ZSCL [50] 4.67 2.35 2.13 2.97 3.15 4.28 4.89 4.70 3.64 MoE-Adapters [48] 2.74 4.71 4.28 1.15 1.50 1.60 2.94 2.77 2.71 Ours 1.70 1.16 0.89 1.04 0.59 1.34 1.12 1.79 1.20 Zero-shot degradation ($\downarrow$) Continual FT 24.81 23.58 19.54 16.46 22.22 19.02 19.54 24.02 21.15 LwF [24] 10.75 10.23 8.63 8.25 12.02 10.33 8.98 11.01 10.03 iCaRL [35] 13.77 12.68 11.28 12.14 13.20 13.20 13.09 14.01 12.92 ZSCL [50] 3.44 3.94 4.02 2.85 3.79 2.31 1.86 1.84 3.00 MoE-Adapters [48] 1.62 2.58 1.04 2.37 4.31 3.05 1.77 0.63 2.17 Ours 1.55 2.04 1.21 1.92 2.79 2.18 1.90 2.08 1.96 Average accuracy ($\uparrow$) Continual FT 76.16 76.24 78.03 68.69 76.64 75.44 72.71 77.45 75.17 LwF [24] 76.78 80.45 80.65 77.52 79.64 79.45 77.31 78.70 78.81 iCaRL [35] 77.99 79.77 79.93 76.66 79.26 79.08 77.06 78.61 78.55 ZSCL [50] 81.89 83.98 84.30 83.49 83.41 82.38 81.92 81.97 82.92 MoE-Adapters [48] 82.71 80.74 81.15 83.97 83.68 83.68 82.73 79.68 82.29 Ours 84.48 84.92 84.97 84.89 85.50 85.07 85.02 84.52 84.92
@@
Method /// Sequence $\mathcal{S}^{1}$ $\mathcal{S}^{2}$ $\mathcal{S}^{3}$ $\mathcal{S}^{4}$ $\mathcal{S}^{5}$ $\mathcal{S}^{6}$ $\mathcal{S}^{7}$ $\mathcal{S}^{8}$ Mean Catastrophic forgetting ($\downarrow$) Continual FT 11.17 10.89 10.16 20.12 10.57 12.14 15.62 9.80 12.56 LwF [24] 9.56 6.38 6.93 11.09 8.37 7.69 10.24 8.44 8.59 iCaRL [35] 8.43 6.90 6.83 10.69 7.09 7.37 10.17 8.66 8.27 ZSCL [50] 4.21 1.41 2.08 3.32 2.85 4.39 5.22 5.13 3.58 Ours 1.92 1.53 0.97 1.14 0.58 1.55 1.29 1.81 1.35 Zero-shot degradation ($\downarrow$) Continual FT 24.54 24.10 19.53 17.60 21.96 18.92 20.26 24.31 21.40 LwF [24] 11.94 11.82 8.27 9.99 13.36 11.47 10.95 12.63 11.30 iCaRL [35] 13.02 12.78 10.89 11.81 12.74 12.87 11.92 13.34 12.42 ZSCL [50] 3.59 4.71 4.17 2.81 3.55 1.97 1.47 2.30 3.07 Ours 1.44 1.80 1.01 1.53 2.17 1.80 1.65 1.82 1.65 Average accuracy ($\uparrow$) Continual FT 75.17 75.13 76.01 67.17 75.54 74.47 71.66 76.40 73.94 LwF [24] 74.14 77.44 77.94 74.89 77.30 77.43 75.50 76.51 76.39 iCaRL [35] 76.97 78.82 78.57 75.43 78.08 78.10 75.70 77.52 77.40 ZSCL [50] 80.49 82.54 82.99 82.08 82.17 80.99 80.30 80.09 81.46 Ours 83.35 83.57 83.88 83.70 84.46 83.82 83.89 83.43 83.76
@@
As shown in Tab. 1, we present the quantitative comparisons with different methods on the MTIL benchmark [50]. The results demonstrate that our method outperforms SOTA CL approaches, showing the effectiveness of our proposed method. By leveraging dual teachers with the proposed teacher selective mechanism, our framework is able to alleviate catastrophic forgetting on all training sequences with less than $2\%$ of performance degradation and properly preserves the zero-shot classification capability.
@@ 4.5 Analysis
The performance on the first dataset in a training sequence inevitably suffers from the most severe catastrophic forgetting during continual learning. In Fig. 4, we plot the evaluation results on the first task through all training rounds to visualize the degree of catastrophic forgetting. The results clearly demonstrate that our method effectively maintains stable performance on the initial task even after multiple training rounds, verifying the effectiveness of our method in alleviating catastrophic forgetting on previous tasks. More experimental results are provided in the supplementary.
@@
In addition to maintaining the knowledge learned from previous tasks, it is also crucial to preserve the zero-shot transferability for continual learning on VLMs. Fig. 5 indicates the level of zero-shot degradation for the last dataset, which experiences the most significant zero-shot degradation. The results show that we are able to keep almost the same zero-shot performance compared with the original pre-trained CLIP, demonstrating the effectiveness of our method in preserving the pre-trained zero-shot classification capability during sequential fine-tuning. More examples on different training sequences are shown in the supplementary.
@@
In Sec. 3.2, we argue that the dual-teacher discrepancy is highly related to determining whether an image is visually similar to previously fine-tuned data. Here, we verify the idea with empirical analysis. At stage $k=2$, given a model $g_{1}$ fine-tuned on the Aircraft dataset ($\mathcal{T}^{1}$) and a pre-trained model $g_{0}$, we calculate the average dual-teacher discrepancy between $g_{1}$ and $g_{0}$ across various datasets, as shown in Tab. 3. The results show that the discrepancy $d$ on the previously fine-tuned data (Aircraft, $\mathcal{T}^{1}$) is significantly greater than on other datasets not fine-tuned by $g_{1}$, while data not aligned with $\mathcal{X}^{1}$ shows a lower discrepancy, empirically supporting our intuition in Eq. 1.
@@
Table 4: An ablation study on knowledge distillation from different teacher selections in $\mathcal{S}^{1}$. Distilling from either $g_{0}$ or $g_{k-1}$ leads to unsatisfactory continual learning and zero-shot performance, respectively. Our method, distilling from both teachers, preserves the zero-shot capability and mitigates catastrophic forgetting in previous tasks.
Method Forgetting ($\downarrow$) Degradation ($\downarrow$) Avg. Accuracy ($\uparrow$) Distill from $g_{0}$ 5.26 2.51 81.35 Distill from $g_{k-1}$ 2.63 3.36 83.61 Ours 1.70 1.55 84.48
4.5.4 Ablation study to the choice of different teachers.
To verify the performance improvement of our proposed dual-teacher distillation mechanism, we conduct an ablation study to different choices of teacher models and present the results in Tab. 4. It can be seen that while distilling from the pre-trained model ($g_{0}$) results in satisfactory zero-shot performance with a drop of $2.51$, it fails to prevent catastrophic forgetting. Conversely, distilling from the most recent model $g_{k-1}$ preserve the continual learning performance but compromise zero-shot capability. Our method, an adaptive distillation scheme, is shown to alleviate catastrophic forgetting while preserving zero-shot performance.
@@
  • After the model has been sequentially trained on multiple rounds (with the last task as Cars), our scheme is still able to select reference images closer to prior tasks (e.g., Aircraft), demonstrating the ability to preserve the earliest fine-tuned knowledge even after multiple downstream tasks, as visualized in the right part of Fig. 6.
@@ 5 Conclusion
In this paper, we propose a Selective Dual-Teacher Knowledge Transfer framework for continual learning that tackles catastrophic forgetting and preserves zero-shot generalization ability simultaneously. By leveraging the most recent fine-tuned and original pre-trained VLMs as dual teachers, our framework selectively distills knowledge based on a dual-teacher discrepancy observed from an auxiliary reference dataset, without requiring label supervision. Comprehensive experiments, including comparisons with state-of-the-art continual learning methods and extensive analysis, quantitatively and qualitatively verify the effectiveness of our framework over existing approaches.
@@
5.0.2 Acknowledgement.
This work is supported in part by the National Science and Technology Council via grant NSTC 112-2634-F-002-007, NSTC 113-2640-E-002-003 and the Featured Area Research Center Program within the framework of the Higher Education Sprout Project by the Ministry of Education 113L900902. We also thank the National Center for High-performance Computing (NCHC) for providing computational and storage resources.
@@ Appendix 0.D Experiments Details
Algorithm 1 Selective Dual-Teacher Knowledge Transfer
Output: The final fine-tuned model $g_{K}$.
1:  for $k$ in $1:K$ do
2:     Freeze $g_{0}$ as the pre-trained knowledge teacher.
3:     Freeze $g_{k-1}$ as the previously fine-tuned knowledge teacher.
4:     Initialize the current model $g_{k}$ by $g_{k-1}$.
5:     for $e$ in $E$ do
6:        while not traverse over all current data $\mathcal{T}^{k}$ do
7:           Sample a batch of current data $B^{k}$.
8:           Sample a batch of ref data $B^{\text{ref}}$.
9:           Calculate $\mathcal{L}_{\text{CE}}$ with the current data $B^{k}$.
10:           Calculate Eq. 3 with $g_{0}$, $g_{k-1}$, and $B^{\text{ref}}$.
11:           Update $g_{k}$ with loss function Eq. 4.
12:        end while
13:     end for
14:  end for
Appendix 0.E The Training Algorithm of Our Proposed Framework
As discussed in Sec. 3.3.1, we provide the detailed training algorithm of our Selective Dual-Teacher Knowledge Transfer framework in Algorithm 1.